# PhoenixOS: Concurrent OS-level GPU Checkpoint and Restore with Validated Speculation

### v2

@@ Abstract
To ensure correct concurrent C/R, PhOS proactively detects GPU memory reads and writes through a two-step process: first, it speculates about GPU memory accesses based on the arguments used when launching GPU kernels; then, it validates these accesses efficiently at runtime using binary instrumentation. With this validated speculation, PhOS retrofits CPU-based concurrent C/R for GPUs through software-based approaches, including soft copy-on-write, soft recopy, and soft on-demand restore. PhOS further proposes several GPU-aware techniques for efficient GPU C/R, including coordinated checkpoint data transfer and execution context pool. For downstream tasks that use C/R for tolerating failures, migrating processes live, and accelerating cold starts in serverless computing, PhOS achieves orders of magnitude higher performance than state-of-the-art OS-level GPU C/R systems like NVIDIA cuda-checkpoint.
@@ Introduction
What is and why concurrent OS-level checkpoint and restore (C/R).   OS-level checkpoint snapshots the execution of a running process as an image, which the OS can then use to restore the execution of the process. It is a foundational OS primitive: cluster management systems rely on it to migrate tenant jobs [63, 74, 55], by first checkpointing the image to the target machine, then restoring from it. Serverless systems leverage OS-level restore to quickly start new processes [15, 32, 20]. Cloud providers periodically checkpoint computing jobs to provide fault tolerance for them [63, 73].
@@
A key feature of OS-level checkpoint and restore is its transparency to the process, making it irreplaceable for functionality and efficiency in various scenarios. For functionality, consider cloud providers that need to migrate tenant jobs to improve cluster utilization and for other purposes [63, 74, 55]. Since these jobs operate as black boxes to the provider [55], migration can only be performed at the OS level. For efficiency, consider serverless platforms that need to quickly start GPU processes [16]. These processes have PyTorch runtime states (e.g., compiler cache) that are tightly intertwined with OS information. Recreating these states from scratch takes seconds; thankfully, OS-level C/R can significantly reduce this overhead by first checkpointing the process state and then restoring it within 10 milliseconds [20, 72, 67].
@@
Additionally, OS-level C/R can also provide superior usability [63, 45], which is why major cloud providers (e.g., Microsoft Forge [45, 28]) employ it for fault tolerance. The key rationale is that performing efficient and correct fault tolerance through user-level checkpointing is challenging due to the broad optimization space and quickly evolving workloads [71, 46, 68]. For instance, Megatron [50], the leading LLM training framework, has recently integrated concurrent checkpoint optimizations proposed three years ago [46], with the checkpoint code now comprising a quarter of its codebase. Meanwhile, many emerging training frameworks, such as those for reinforcement learning [57], still lack these optimizations. With efficient OS-level checkpointing, applications can avoid the burden of implementing and optimizing their own fault tolerance mechanisms.
Concurrent OS-level C/R—performing checkpoint and restore while allowing processes to run concurrently—is becoming increasingly important, as stopping GPU processes severely impacts application performance due to the lengthy data copy time during the C/R process (§2.3). For example, Microsoft reports that over 3.9% of GPU users experience quality issues from process migration stalls using C/R [26]. The stall caused by restoring a Llama-2-13B inference using state-of-the-art OS-level C/R [63] requires a 6.2-seconds pause—31$\times$ longer than the Time-To-First-Token (TTFT) for inference, which significantly hinders deploying GPU processes in emerging serverless computing [33, 16]. Finally, in fault tolerance scenarios like model training, the checkpoint time can be comparable to the iteration time (46–87%, see §8), and hiding them with concurrent execution can save precious GPU hours [71, 68, 46, 31].
@@
Current systems and the key challenge.   Existing systems like NVIDIA’s cuda-checkpoint [63, 56] cannot perform concurrent checkpoint and restore of GPU-related states. While concurrent execution of C/R is standard practice for CPU processes, it is ineffective when the GPU is stopped because the CPU must wait for GPU execution results (§2.3). This raises a key question: Can we achieve efficient and effective OS-level checkpoint and restore during concurrent GPU execution?
@@
The key challenge is ensuring correctness—a checkpoint must reflect a valid process state as if no checkpoint had occurred (§4). For example, if the GPU executes two kernels before the checkpoint, the checkpoint must capture the exact state of GPU memory after these two kernels write. However, the process running concurrently can overwrite data that has not been checkpointed, resulting in an incorrect checkpoint. Similarly, during the restore, if a GPU kernel reads data while the OS is concurrently restoring it, the kernel may read incorrect data, corrupting the process state.
@@
Table 1: An overview of the information (Info.) required for correct concurrent checkpoint and restore, and how PhOS tracks them.
@@
Checkpoint Writeset Permission [73, 66] and dirty bits [14] Speculation + Validation
Restore Readset Present bits [72, 20, 69] 4.1)
@@
The key for correctness is to trace the read and write sets of the concurrent execution, i.e., which bytes of CPU/GPU states are read and written by the process. Take checkpointing as an example: the OS can use this information to isolate the writes that cause an incorrect checkpoint with copy-on-write [61, 73], or recopy the writes to the checkpoint [14] for correctness. For CPU states, the OS can leverage OS-mediated data paths with hardware paging to gather this information, as summarized in Table 1. However, GPUs lack the necessary hardware support [27] and bypass the OS during their execution to maximize performance.
@@
Key insights.   Unlike CPU, whose execution is a coarse-grained black box to the OS, the execution of GPU is composed of fine-grained units (e.g., kernels), whose control flow is mediated by the OS. Specifically, applications trigger GPU execution through fine-grained GPU API calls (e.g., CUDA [49]), which the OS can intercept to trace their read and write sets for concurrent checkpoint and restore. Because each API will trigger fine-grained GPU state modifications (§8.4), such an API-level tracing can realize an efficient and correct concurrent C/R.
@@
However, tracing read and write sets at the API level is non-trivial, because although some APIs have well-defined read and write semantics, e.g., those launch kernels provided by the vendors like cuBLAS [52], the process can also trigger user-developed kernels that may contain arbitrary code. Fortunately, we find that although such kernels may include complex code, each typically serves a clear and straightforward computational purpose, such as matrix multiplication or convolution operations. This makes it possible to accurately speculate which data will be accessed by simply analyzing the kernel’s launch arguments (§4.1).
@@
The PhoenixOS (PhOS).  Based on our insights, we built PhOS, the first OS service capable of concurrently checkpointing and restoring GPU processes. PhOS speculates the data accessed by each GPU execution unit using its launch arguments, and uses this speculated information to retrofit CPU-based concurrent C/R protocols for GPUs, including soft copy-on-write (§4.2), soft recopy (§4.3), and soft on-demand restore (§6). To ensure correctness, we further instrument a lightweight validator to correctly handle speculation failures.
@@
Retrofitting CPU-based concurrent C/R techniques on GPUs with our validated speculation—although ensures correctness—still suffers from performance issues if not considering the features of GPUs carefully. First, during concurrent checkpoint, the interference between CPU and GPU checkpoint, as well as between application data transfers and checkpoint transfers, can significantly impact the efficiency and effectiveness of the concurrent checkpoint. To mitigate these interferences, we design a coordinated and prioritized checkpoint mechanism (§5). Second, before allowing a concurrent restore, we must create the proper GPU execution environment, which incurs overhead comparable to that of data restoration itself. We propose using a GPU context pool to bypass the context creation and fully unleash the power of concurrent restore (§6).
@@
Demonstrations.   PhOS runs on Linux and it can C/R unmodified GPU applications (including multi-GPU processes) on NVIDIA GPUs. We chose NVIDIA GPUs due to their prevalent adoption, despite higher implementation complexity. Nevertheless, our design also generalizes to other GPUs and accelerators, because they follow the same execution model PhOS assumed for NVIDIA GPUs (§2.1). On A800 GPU servers, when compared to state-of-the-art systems—Singularity [63] and NVIDIA’s official OS-level C/R (cuda-checkpoint) [56], PhOS reduces the application stall caused by checkpoint and restore by up to 160%. More importantly, PhOS brings orders of magnitude improvements on end-to-end application performance across various tasks. Notably, compared to Singularity, it reduces the wasted GPU time by 76% when training Llama3-13B on multiple GPUs, it reduces the downtime caused by migrating a Llama2-13B inference job from 9.8 to 2.3 seconds, and it can cold-start a Llama2-13B inference job within only 622 milliseconds, which is 124–450% and 114–342% faster than cuda-checkpoint and Singularity, respectively.
@@
  • The first efficient GPU execution read and write set detection via validated speculation (§4.1).
  • The first set of OS-level concurrent checkpoint and restore protocols for GPUs (§4.2, §4.3, and §6) that ensure correctness with validated speculation.
@@
PhOS will be open-sourced upon publication.
@@
Figure 1: Illustrations of (a) how a GPU process executes, (b) how a stop-the-world OS-level checkpoint and restore works, (c) how PhOS does checkpoint and restore, and (d) the main content of checkpointed data.
GPUs are accelerators with massive multithreading typically attached to the CPU via PCIe. Programs executed on GPUs are termed kernels, which contain machine code (e.g., SASS [54]) that is either pre-compiled or just-in-time compiled. Kernels process data residing in GPU buffers, where each buffer constitutes a contiguous GPU virtual memory region with application-controlled granularity.
When a process starts, the GPU driver creates proper execution context, e.g., CUcontext [53], involving loading or compiling the kernel binaries and configuring the GPU virtual memory [42]. Afterward, processes use standardized GPU driver or toolkit APIs (we term GPU API in this paper) like CUDA [49], illustrated in Figure 1 (a), to trigger computations on GPUs. For example, the process can launch kernels via cudaLaunchKernel and copy CPU data to GPU using cudaMemcpy. Due to the limited control capability of GPUs, the GPU execution is orchestrated by the CPU by calling GPU APIs based on the application’s logic. PhOS intercepts GPU APIs for realizing checkpointing and restoration, thus supporting all GPU applications at the OS-level.
@@
Basic checkpoint and restore. A checkpoint captures a process’s execution state at a specific time in an image (termed as checkpoint). The OS can then use this checkpoint to recreate (restore) the process. A checkpoint includes two data categories: data state (data in the CPU virtual memory and GPU buffers) and control state (e.g., CPU registers that store the program counter, PC). The checkpointed CPU states also include kernel objects like network connections [17, 66].
A common approach for implementing GPU checkpointing is by quiescing process execution (both CPU and GPU), i.e., stops the CPU execution and waiting for all in-flight GPU kernels and communications to finish, then copying all aforementioned data to the checkpoint [56, 63]. Figure 1 (b) illustrates this. For restoration, the OS first loads data state into CPU memory and GPU buffers, then restores control state to resume execution.
Concurrent CPU checkpoint and restore. To prevent application stall during checkpointing and restoration, which has non-trivial overhead, OS researchers have investigated concurrent CPU checkpointing and restoration for decades [73, 66, 14, 72, 20, 69]. Concurrent checkpointing allows CPU execution during data copying, thus hiding the overhead of checkpointing. For correctness, the OS either isolates concurrent CPU writes via copy-on-write [73, 66], or recopies concurrently written data to the checkpoint [14]. Concurrent restoration enables immediate process resumption, i.e., no need to wait for the restoration to finish. During the process execution, the data is being concurrently copied from the checkpoint to the CPU memory. For correctness, if the process touches non-restored data during execution, the CPU will trigger a page fault, so the OS will copy the missing data from the checkpoint on demand [72, 20, 69].
@@
Figure 2: A breakdown of checkpoint and restore overhead.
Figure 2 analyzes the stalls caused by checkpoint and restore on a Llama2-13B inference process. We evaluate Singularity [63]—the state-of-the-art GPU checkpoint and restore system, other baselines like cuda-checkpoint [56] are much slower. We have enabled concurrent CPU checkpointing and restoration using CRIU’s incremental dump and restore [19]. The detailed setup and performance of cuda-checkpoint can be found in §8.
Copying GPU data. Because the GPU is stopped when copying data to and from checkpoint, the applications are stalled despite the CPU-side of the process being concurrently executed. The stall time is non-trivial: checkpointing and restoring data each require more than 2.1 seconds respectively when transferred with PCIe, causing thousands of tokens to be disrupted. The stall further scales with the GPU memory used, which is substantial: For instance, Llama2-13B inference occupies 56 GB of active memory. Given a typical 32 GBps PCIe 4.0 CPU–GPU link, OS-level checkpoint and restore takes at least 1.75 seconds.
Creating GPU contexts. Before restore can copy data to the GPU, the OS must create proper GPU contexts. This establishes a restoration barrier as context creation incurs comparable overheads to data copying, i.e., 3.1 vs. 1.7 seconds in our motivation experiment. GPU context initialization exceeds CPU context creation time due to complex hardware configuration and driver state loading [42]. In comparison, creating CPU context, i.e., a Linux process, merely initializes a small number of kernel data structures.
The PhOS Operating System
Design goals.   Our goals are efficient and correct OS-level GPU checkpoint and restore. For efficiency, the goal is to minimize process stalls during checkpoint or restore. PhOS achieves this through: concurrent process execution during data copying for both checkpoint and restore, and bypassing GPU context creation during restore, as illustrated in Figure 1 (c). For correctness, the goal is to make the checkpoint reflect an application state indistinguishable from a non-checkpointed execution [63, 37, 66].
Figure 3: PhOS system architecture. Note that IPC and LFC stand for inter-process call and local function call, respectively.
System architecture and execution flow. Figure 3 shows PhOS’s architecture that has three core system components: a command-line tool (❶), a per-process frontend library to facilitate C/R (❷), and a backend module to do the C/R (❸).
@@
Once a checkpoint or restore command is received by the tool, it will communicate with the frontend library (❷) embedded in the GPU driver and its toolkits. The frontend will call our backend to do the checkpoint and restore, as well as trace the read and write sets (§4.1) when the process calls a GPU API. This information is provided to our C/R backend to ensure correctness (§4 and §6). Our current implementation leverages LD_PRELOAD to extend the CUDA GPU API as CUDA is closed-source, but LD_PRELOAD is not a requirement once the code is available.
Our backend (❹) uses CRIU [18] to checkpoint and restore CPU process states, and our efficient GPU engine (see §5) for GPU states. Choosing CRIU allows PhOS to correctly checkpoint and restore complex states like network connections. We support a wide range of checkpoint media: PhOS can read and write checkpoints to local SSD, CPU DRAM and even the DRAM of another machine via RDMA (§7), which is critical for efficient process migration. The backend also includes a GPU context pool—containing pre-allocated contexts that avoid creating GPU contexts during restore (§6).
PhOS command-line tool uses inter-process calls (IPC) to communicate with the frontend. On the other hand, the frontend and backend can communicate either via IPC or local function calls (LFC). We prefer LFC when feasible by linking the frontend and backend in the same library. However, our context pool requires storing the pool in an OS service daemon, so LFC is not possible. Thus, for scenarios that require the context pool, e.g., serverless startup [33, 16], we use IPC. For others like fault tolerance we employ LFC. The IPC overhead remains minimal because GPU APIs execute asynchronously, allowing overlapping IPC cost with GPU execution [70, 76].
@@
Correctness guarantee of PhOS. Intuitively, a checkpoint is correct if it could occur in a checkpoint-free execution [23]. A stop-the-world checkpoint described in §2.2 naturally guarantees correctness. PhOS ensures correctness by ensuring that our checkpoint matches the one that may come from the existing stop-the-world checkpoint.
Figure 4: An illustration of correctness guarantee of concurrent checkpointing enabled by different protocols of PhOS.
Overview of the checkpoint protocols. We retrofit two CPU-inspired protocols [73, 66, 14]. As Figure 4 shows, consider our checkpoint starting at t1 with concurrent data copy done at t2. Compared with a stop checkpoint at t1, inconsistencies are caused by new writes issued between [t1,t2]. Our soft copy-on-write (CoW) protocol isolates these writes (§4.2), making the checkpoint only copy a frozen GPU state at t1. Alternatively, compared to a stop checkpoint at t2, our approach may miss new writes. Thus, at t2, our soft recopy protocol first quiesces the process and recopies the missed writes to ensure correctness (§4.3). CoW is commonly used in fault tolerance cases, where the application can tolerate restoring from a stale state (t1). On the other hand, recopy is used in cases where the restored process must be resumed from the latest execution (t2), e.g., live migration (§7).
Both protocols require tracing the process’s write set, i.e., which bytes are written during concurrent checkpoint. For CPU writes, we follow the traditional approach to use permission bits in the page table, i.e., write protection for copy-on-write and dirty bits to trace at the page granularity. We omit the detailed CPU-side description since it is well studied. Unfortunately, modern GPUs lack these support, e.g., no hardware dirty bits [27]. Therefore, before delving into our protocols, we first propose a software-based approach to trace GPU writes.

### v3

@@ Abstract.
To ensure correct concurrent C/R, PhOS proactively detects GPU memory reads and writes through a two-step process: first, it speculates about GPU memory accesses based on the arguments used when launching GPU kernels; then, it validates these accesses efficiently at runtime using binary instrumentation. With this validated speculation, PhOS retrofits CPU-based concurrent C/R for GPUs through software-based approaches, including soft copy-on-write, soft recopy, and soft on-demand restore. PhOS further proposes several GPU-aware techniques for efficient GPU C/R, including coordinated checkpoint data transfer and execution context pool. For downstream tasks that use C/R for tolerating failures, migrating processes between machines, and accelerating cold starts in serverless computing, PhOS achieves orders of magnitude higher performance than state-of-the-art OS-level GPU C/R systems like NVIDIA cuda-checkpoint.
[…]
@@ Introduction
What is and why concurrent OS-level checkpoint and restore (C/R).    OS-level checkpoint snapshots the execution of a running process as an image, which the OS can then use to restore the execution of the process. It is a foundational OS primitive: cluster management systems rely on it to migrate tenant jobs (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/osdi/XiaoBRSKHPPZZYZ18, ; cuda-migration, )—by first checkpointing the image to the target machine, then restoring from it. Serverless systems leverage OS-level restore to quickly start new processes (serverless-gpu-ali, ; aws-snapstart, ; DBLP:conf/asplos/DuYXZYQWC20, ). Cloud providers periodically checkpoint computing jobs to provide fault tolerance in case of failures (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/sosp/Wu0M023, ).
@@
A key feature of OS-level checkpoint and restore is that it is transparency to the processes, making it irreplaceable for functionality and efficiency. For functionality, consider cloud providers that need to migrate tenant jobs to improve cluster utilization and for other purposes (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/osdi/XiaoBRSKHPPZZYZ18, ; cuda-migration, ). Since these jobs operate as black boxes to the providers (cuda-migration, ), migration can only be performed at the OS level. For efficiency, consider serverless platforms that need to quickly start GPU processes (ali-serverless-gpu, ). These processes have PyTorch runtime states (e.g., compiler cache) that are tightly intertwined with OS information. Recreating these states from scratch takes seconds; thankfully, OS-level C/R can significantly reduce this overhead to milliseconds (DBLP:conf/asplos/DuYXZYQWC20, ; DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/asplos/Ustiugov0KBG21, ).
@@
Additionally, OS-level C/R can also provide superior usability (DBLP:journals/corr/abs-2202-07848, ; microsoft-forge, ), which is why major cloud providers (e.g., Microsoft Forge (microsoft-forge, ; DBLP:conf/eurosys/GuptaKKVGKRS24, )) employ it for fault tolerance. The key rationale is that performing efficient and correct fault tolerance through user-level C/R is challenging due to the broad optimization space and quickly evolving workloads (DBLP:conf/sosp/WangJZZFNW23, ; DBLP:conf/fast/MohanPC21, ; DBLP:journals/corr/abs-2407-20143, ). For instance, the leading LLM training framework Megatron (megatron-checkpoint, ) has recently integrated concurrent checkpoint optimizations proposed three years ago (DBLP:conf/fast/MohanPC21, ), with the checkpoint code now comprising a quarter of its codebase. Meanwhile, many emerging training frameworks, such as those for reinforcement learning (gym, ), still lack these features. OS-level C/R avoids the developers’ burden of implementing and optimizing their own fault tolerance mechanisms.
[…]
@@
Current systems and the key challenge.    Existing systems like NVIDIA cuda-checkpoint (DBLP:journals/corr/abs-2202-07848, ; nvidia-new-ckpt-driver-repo, ) cannot perform concurrent checkpoint and restore of GPU-related states. While concurrent execution of C/R is standard practice for CPU processes, it is ineffective when the GPU is stopped because the CPU must wait for GPU execution results (§2.3). This raises a key question: Can we achieve efficient and effective OS-level checkpoint and restore during concurrent GPU execution?
@@
The key challenge is ensuring correctness—a checkpoint must reflect a valid process state as if no checkpoint had occurred (§4). For example, if the GPU executes two kernels before the checkpoint, the checkpoint must capture the exact state of GPU memory after these two kernels write. However, the process running concurrently can corrupt the checkpoint by overwriting data that has not been checkpointed. Similarly, during the restore, if a GPU kernel reads data while the OS is concurrently restoring it, the kernel may read incorrect data, corrupting the process state.
@@
Table 1. An overview of the information (Info.) required for correct concurrent checkpoint and restore, and how PhOS traces them.
@@
Checkpoint Writeset Permission (DBLP:conf/sosp/Wu0M023, ; Aurora, ) and dirty bits (DBLP:conf/nsdi/ClarkFHHJLPW05, ) Speculation + Validation
Restore Readset Present bits (DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/asplos/DuYXZYQWC20, ; DBLP:conf/eurosys/WangHW19, ) 4.1)
@@
The key for correctness is to trace the read and write sets of the concurrent execution, i.e., which bytes of CPU/GPU states are read and written by the process. Take checkpointing as an example: the OS can use this information to isolate the writes that cause an incorrect checkpoint with copy-on-write (DBLP:conf/sosp/ShapiroSF99, ; DBLP:conf/sosp/Wu0M023, ), or recopy the writes to the checkpoint (DBLP:conf/nsdi/ClarkFHHJLPW05, ) for correctness. For CPU states, the OS can leverage OS-mediated data paths with hardware paging to gather this information, as summarized in Table 1. However, GPUs lack the necessary hardware support (DBLP:conf/uss/0002Z024, ) and bypass the OS during their execution to maximize performance.
@@
Key insights.    Unlike CPU, whose execution is an entire black box to the OS, the execution of GPU is composed of fine-grained units (e.g., kernels), whose control flow is mediated by the OS. Specifically, applications trigger GPU execution through fine-grained GPU API calls (e.g., CUDA (cuda-driver, )). Moreover, each API will trigger fine-grained GPU state modifications (§8.5), so the OS can intercept these APIs to trace their read and write sets for concurrent checkpoint and restore.
@@
However, tracing read and write sets at the API level is non-trivial, because although some APIs have well-defined read and write semantics, e.g., those launch kernels provided by the vendors like cuBLAS (cuBLAS, ), the process can trigger user-developed kernels with arbitrary code. Fortunately, we find that although such kernels may include complex code, each typically serves a clear computational purpose, such as doing computations on a set of GPU buffers. This makes it possible to accurately speculate which data will be accessed by simply analyzing the kernel’s launch arguments (§4.1).
@@
The PhoenixOS (PhOS).    Based on our insights, we built PhOS, the first OS service capable of concurrently checkpointing and restoring GPU processes (DBLP:journals/corr/abs-2405-12079, ). PhOS speculates the data accessed by each GPU kernel using its launch arguments, and uses this speculated information to retrofit CPU-based concurrent C/R protocols for GPUs, including copy-on-write (§4.2), recopy (§4.3), and on-demand restore (§6). To ensure correctness, we further instrument a lightweight validator to correctly handle speculation failures.
@@
Retrofitting CPU-based concurrent C/R techniques on GPUs with our validated speculation—although ensuring correctness—still suffers from performance issues if not considering GPU features carefully. First, during concurrent checkpoint, the interference between CPU and GPU checkpoint, as well as between application data transfers and checkpoint transfers, can significantly impact the efficiency of the concurrent checkpoint. To mitigate these interferences, we design a coordinated and prioritized checkpoint mechanism (§5). Second, before allowing a concurrent restore, we must create the proper GPU execution environment, which incurs overhead comparable to that of data restoration itself. We propose using a GPU context pool to bypass the context creation to fully unleash the power of concurrent restore (§6).
Figure 1. Illustrations of (a) how a GPU process executes, (b) how a stop-the-world OS-level checkpoint and restore works, (c) how PhOS does concurrent checkpoint and restore, and (d) the main content of checkpointed data.
@@
Demonstrations.  PhOS can C/R unmodified GPU applications (including multi-GPU processes) on NVIDIA GPUs. We chose NVIDIA GPUs due to their prevalent adoption, despite higher implementation complexity. Nevertheless, our design also generalizes to other GPUs and accelerators, because they follow the same execution model of NVIDIA GPUs (§2.1). On A800 GPU servers, when compared to state-of-the-art systems—Singularity (DBLP:journals/corr/abs-2202-07848, ) and NVIDIA’s official OS-level C/R (cuda-checkpoint) (nvidia-new-ckpt-driver-repo, ), PhOS reduces the application stall caused by checkpoint and restore by up to 160%. More importantly, PhOS brings orders of magnitude end-to-end application performance improvements. Notably, compared to Singularity, it reduces the wasted GPU time caused by fault tolerance by 76% when training Llama3-13B on multiple GPUs, it reduces the downtime caused by migrating a Llama2-13B inference job from 9.8 to 2.3 seconds, and it can cold-start a Llama2-13B inference job within only 622 milliseconds, which is 124–450% and 114–342% faster than cuda-checkpoint and Singularity, respectively.
@@
  • The first efficient GPU execution read and write set tracing method via validated speculation (§4.1).
  • The first set of OS-level concurrent checkpoint and restore protocols for GPUs: §4.2, §4.3, and §6.
@@
PhOS is open-source and publicly available at https://github.com/SJTU-IPADS/PhoenixOS.
@@
GPUs are accelerators with massive multithreading capability typically attached to the CPU via PCIe. Programs executed on GPUs are termed kernels, which contain machine code (e.g., SASS (sass, )) that is either pre-compiled or just-in-time compiled. Kernels process data residing in GPU buffers, where each buffer constitutes a contiguous GPU virtual memory region with application-controlled granularity.
When a process starts, the GPU driver creates execution context, e.g., CUcontext (cuda-context, ), involving compiling and loading the kernel binaries, configuring the GPU virtual memory and others (Mai2023HONEYCOMB, ). Afterward, processes use GPU driver or toolkit APIs (we term GPU API in this paper) like CUDA (cuda-driver, ), illustrated in Figure 1 (a), to trigger computations on GPUs. For example, the process can launch kernels via cudaLaunchKernel and copy CPU data to GPU (and vice versa) using cudaMemcpy. PhOS intercepts GPU APIs for realizing checkpointing and restoration, thus supporting all GPU applications at the OS-level.
@@
Basic checkpoint and restore.    A checkpoint captures a process’s execution state at a specific time in an image (termed as checkpoint). The OS can then use this checkpoint to recreate (restore) the process. A checkpoint includes two data categories: data state (data in the CPU virtual memory and GPU buffers) and control state (e.g., CPU registers that store the program counter, PC). The checkpointed CPU states also include kernel objects like network connections (criu-tcp, ; Aurora, ).
A common approach for implementing GPU checkpointing is by first quiescing process execution (both CPU and GPU), i.e., stops the CPU execution and waiting for all in-flight GPU kernels and communications to finish, then copying all aforementioned data to the checkpoint (nvidia-new-ckpt-driver-repo, ; DBLP:journals/corr/abs-2202-07848, ). Figure 1 (b) illustrates this. For restoration, the OS first loads data state into CPU memory and GPU buffers, then restores control state to resume execution.
Concurrent CPU checkpoint and restore.    To prevent application stall during checkpointing and restoration, which has non-trivial overhead, OS researchers have investigated concurrent CPU checkpointing and restoration for decades (DBLP:conf/sosp/Wu0M023, ; Aurora, ; DBLP:conf/nsdi/ClarkFHHJLPW05, ; DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/asplos/DuYXZYQWC20, ; DBLP:conf/eurosys/WangHW19, ). Concurrent checkpointing allows CPU execution during data copying, thus hiding the overhead of checkpointing. For correctness, the OS either isolates concurrent CPU writes via copy-on-write (DBLP:conf/sosp/Wu0M023, ; Aurora, ), or recopies concurrently written data to the checkpoint (DBLP:conf/nsdi/ClarkFHHJLPW05, ). Concurrent restoration enables immediate process resumption, i.e., no need to wait for the data to be fully restored. During the process execution, the data is being concurrently copied from the checkpoint to the CPU memory. For correctness, if the process touches non-restored data, the CPU will trigger a page fault, so the OS will copy the missing data from the checkpoint on demand (DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/asplos/DuYXZYQWC20, ; DBLP:conf/eurosys/WangHW19, ).
@@
Figure 2. A breakdown of checkpoint and restore overhead.
Figure 2 analyzes the stalls caused by checkpoint and restore on a Llama2-13B inference process. We evaluate Singularity (DBLP:journals/corr/abs-2202-07848, )—the state-of-the-art GPU checkpoint and restore system, other baselines like cuda-checkpoint (nvidia-new-ckpt-driver-repo, ) are much slower. We have enabled concurrent CPU checkpointing and restoration using CRIU’s incremental dump and restore (criu-memory-change-tracking, ). The detailed setup and performance of cuda-checkpoint can be found in §8.
Copying GPU data.    Because the GPU is stopped when copying data to and from checkpoint, the applications are stalled despite the CPU-side of the process being concurrently executed. The stall time is non-trivial: checkpointing and restoring data each require more than 2.1 seconds respectively when transferred with PCIe, causing thousands of tokens to be disrupted. The stall further scales with the GPU memory used, which is substantial: For instance, Llama2-13B inference occupies 55 GB of active memory. Given a typical 32 GBps PCIe 4.0 CPU–GPU link, OS-level checkpoint and restore takes at least 1.7 seconds.
Creating GPU contexts.    Before restore can copy data to the GPU, the OS must create proper GPU contexts. This establishes a restoration barrier as context creation incurs comparable overheads to data copying, i.e., 3.1 vs. 1.7 seconds in our motivation experiment. GPU context initialization exceeds CPU context creation time due to complex hardware configuration and driver state loading (Mai2023HONEYCOMB, ). In comparison, creating CPU context, i.e., a Linux process, merely initializes a small number of kernel data structures.
The PhoenixOS (PhOS)
Design goals.   Our goals are efficient and correct OS-level GPU checkpoint and restore. For efficiency, the goal is to minimize process stalls during checkpoint or restore. PhOS achieves this through: concurrent process execution during data copying for both checkpoint and restore, and bypassing GPU context creation during restore, as illustrated in Figure 1 (c). For correctness, the goal is to make the checkpoint reflect an application state indistinguishable from a non-checkpointed execution (DBLP:journals/corr/abs-2202-07848, ; DBLP:conf/ics/LeeSHTKE19, ; Aurora, ).
Figure 3. PhOS system architecture. IPC and LFC stand for inter-process call and local function call, respectively.
System components and execution flow.    Figure 3 shows PhOS’s architecture that has three core system components: a command-line tool (❶), a per-process frontend library to facilitate C/R (❷), and a backend module to do the C/R (❸).
@@
Once a checkpoint or restore command is received by the tool, it will communicate with the frontend library (❷) embedded in the GPU driver and its toolkits. The frontend will further call our backend to do the checkpoint and restore, as well as trace the read and write sets (§4.1) when the process calls a GPU API. This information is provided to our C/R backend to ensure correctness (§4 and §6). Our current implementation leverages LD_PRELOAD to extend the CUDA GPU API as CUDA is closed-source, but LD_PRELOAD is not a requirement once the code is available.
Our backend (❸) uses CRIU (criu, ) to checkpoint and restore CPU process states, and our efficient GPU engine (see §5) for GPU states. Choosing CRIU allows PhOS to correctly checkpoint and restore complex states like network connections. However, CRIU may not be the most efficient solution for CPU checkpointing. Our current implementation overlaps the CRIU checkpointing with our GPU checkpointing but a faster CPU checkpointing would make PhOS easier to implement.
PhOS backend supports a wide range of checkpoint media: it can read and write checkpoints to local SSD, CPU DRAM and even the DRAM of another machine via RDMA (§7), which is critical for efficient process migration. The backend also includes a GPU context pool—containing pre-allocated contexts that avoid creating GPU contexts during restore (§6).
PhOS command-line tool uses inter-process calls (IPC) to communicate with the frontend. On the other hand, the frontend and backend can communicate either via IPC or local function calls (LFC). We prefer LFC whenever feasible by linking the frontend and backend in the same library. However, IPC is required for cases when the applications require accessing the context pool, e.g., for fast restore (see §7). This is because the context must be pre-created in PhOS daemon to bypass the costly context creation during restore (see Figure 2). This comes at the cost of extra GPU API calling overhead during application running (excluding the restore), e.g., we observed a maximum of 9% IPC overhead for typical AI applications after applying recent works in accelerating calling GPU APIs with IPC (DBLP:journals/corr/abs-2401-13354, ; DBLP:journals/corr/abs-2306-03622, ). We believe it is a proper trade-off as the restore time dominates the overall application execution time in the required scenarios (see §7).
@@
Correctness guarantee of PhOS.    Intuitively, a checkpoint is correct if it could occur in a checkpoint-free execution (DBLP:journals/csur/ElnozahyAWJ02, ). A stop-the-world checkpoint described in §2.2 naturally guarantees correctness. PhOS ensures correctness by ensuring that our checkpoint matches the one that may come from the existing stop-the-world checkpoint.
Figure 4. An illustration of correctness guarantee of concurrent checkpointing enabled by different protocols of PhOS.
Overview of the checkpoint protocols.    We retrofit two CPU-inspired protocols (DBLP:conf/sosp/Wu0M023, ; Aurora, ; DBLP:conf/nsdi/ClarkFHHJLPW05, ) for a correct concurrent GPU checkpoint. As Figure 4 shows, consider our checkpoint starting at t1 with concurrent data copy done at t2. Compared with a stop checkpoint at t1, inconsistencies are caused by new writes issued between [t1,t2]. Our soft copy-on-write (CoW) protocol isolates these writes (§4.2), making the checkpoint only copy a frozen GPU state at t1. Alternatively, compared to a stop checkpoint at t2, our approach may miss new writes. Thus, at t2, our soft recopy protocol quiesces the process and recopies the missed writes to ensure correctness (§4.3). CoW is commonly used in fault tolerance cases, where the application can tolerate restoring from a stale state (t1). On the other hand, recopy is used in cases where the restored process must be resumed from the latest execution (t2), e.g., live migration (§7).
Both protocols require tracing the process’s write set, i.e., which bytes are written during concurrent checkpoint. For CPU writes, we follow the traditional approach to use permission bits in the page table, i.e., write protection for copy-on-write and dirty bits to trace at the page granularity. We omit the detailed CPU-side description since it is well studied. Unfortunately, modern GPUs lack these support, e.g., no hardware dirty bits (DBLP:conf/uss/0002Z024, ). Therefore, before delving into our protocols, we first propose a software-based approach to trace GPU writes.